OpenStudy (anonymous):

x^2+y^2+12x+8y=48 Please don't just give me the answer I need to understand the steps... it's making me so frustrated. Help!!

OpenStudy (anonymous):

Alrighty! Just gimme a sec to work out the problem on my own, and then I can help :)

OpenStudy (anonymous):

thanks you a million times!!

OpenStudy (anonymous):

Okay. You can't factor this out normally, sadly, enough. However, you could split up 48 into 36 and 12.

OpenStudy (anonymous):



OpenStudy (anonymous):

And have 12 for your y equation :)

hero (hero):

I assume that you have to write the equation of the circle in standard form.

OpenStudy (anonymous):

okay its a conic equation but i dont know which one but yeah

hero (hero):

All that needs to be done is to complete the square.

OpenStudy (anonymous):

i can graph it but i cant figure out which standard form to put it into

hero (hero):

(x - h)^2 + (y - k)^2 = r^2

OpenStudy (anonymous):

@Hero : I don't know conics all too well. >.< Or at least what they are?

OpenStudy (anonymous):

I didn't realize this was a conic equation until afterwards xD

hero (hero):

Isolate terms with x on one side: x^2 + 12x = 48 - y^2 - 8y Then complete the square by adding 36 to both sides: x^2 + 12x + 36 = 48 + 36 - y^2 - 8y (x + 6)^2 = 84 - y^2 - 8y Isolate terms with y on one side: y^2 + 8y = 84 - (x + 6)^2 Then complete square by adding 16 to both sides: y^2 + 8y + 16 = 84 + 16 - (x + 6)^2 (y + 4)^2 = 100 - (x + 6)^2 Then put the equation in standard form: (x + 6)^2 + (y + 4)^2 = 100 (x + 6)^2 + (y + 4)^2 = 10^2

OpenStudy (anonymous):

Ah, okay. Makes sense. :D

OpenStudy (anonymous):

thank youu!!

hero (hero):

yw